The University of Smallville requires its students to pass an examination in college-level mathematics before they can graduate. The students are given three chances to pass the exam; 61 % pass it on their first attempt; 66 % of those that take it a second time pass it then; and 42 % of those that take it a third time pass it then.
a. What percent of the students pass the exam? %
b. What percent of the students are not allowed to graduate because of their performance on the exam? %
c. What percent of the students take the exam at least twice? %
d. What percent of the students take the test three times? %

Answer by edjones(8007)

61% pass the 1st test
.
100-61=59 fail the 1st
59*.66=24.74% of the total pass the 2nd.
.
100-61-24.74=14.26% of total fail the 2nd
14.26*.42=5.99 of the total pass the 3rd.
.
a)61+24.74+5.99=91.73% pass
.
Ed